# The Art of Storytelling: Evaluating AI Models for Storytelling

### v5

@@ Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling. We evaluate GenAI-driven narrative co-creation, text-to-speech conversion, text-to-music and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Evaluating AI Models for Storytelling
Samee Arif1, Taimoor Arif2, *Muhammad Saad Haroon1, *Aamina Jamal Khan1, Agha Ali Raza1, Awais Athar3 1Lahore University of Management Sciences, 2University of Michigan, 3EMBL European Bioinformatics Institute {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@ebi.ac.uk
@@ 1 Introduction
In this paper, we evaluate a the feasibility of GenAI to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills Ciğerci and Yıldırım (2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.
The system leverages the capabilities of Large Language Models (LLMs) to generate narratives based on user prompts, Text-to-Speech (TTS) models to convert these stories into spoken word, Text-to-Video (TTV) models to visualize the narratives through animation, and a Text-to-Music (TTM) model to generate background music that complements the visuals and enhances the storytelling experience. This combination of technologies creates an educational tool, supporting both creative expression and cognitive development, enhancing the overall learning experience for children.
@@
The conversion of written text into spoken words, providing auditory stimuli that can improve reading comprehension, especially for students with learning disabilities. Studies have shown that TTS can support struggling readers by offering audio input as digital text is read aloud, aiding in comprehension and retention (Keelor et al., 2020). Visual stimuli can enhance understanding and retention of information. Research indicates that combining text and images in multimedia explanations is more effective than text alone, as it caters to diverse learning preferences and helps maintain attention (Liu, 2024). Background music adds an emotional layer to narratives, facilitating deeper engagement and supporting memory formation. Music has been recognized for its ability to evoke emotions and create a captivating environment, which can enhance memory retention by increasing engagement (Sinoradzki, 2023). By integrating these components, our system addresses the limitations of traditional storytelling methods that often lack multimodal stimuli, leading to reduced attention spans and engagement. This multimodal approach ensures that storytelling is not only entertaining but also educational, catering to various learning styles and promoting cognitive development.
Our contributions can be summarized as follows:
  1. 1.
    Evaluation Framework: We present an evaluation framework and conduct human evaluation of the system components. Our evaluation includes assessing LLMs for story generation and filtering out content inappropriate for children, as well as the TTS and TTV models for their respective roles in the storytelling process. This evaluation ensures that the system delivers content that is both engaging and suitable for young audiences.
  2. 2.
    Public Datasets: We publicly release the human evaluation data, that can be used as a benchmark for assessing LLMs as evaluators in story generation and content review tasks. Additionally, we provide two benchmarking datasets for filtering inappropriate content for children, promoting research on child-safe AI systems.
The code, model outputs and the evaluation scores are publicly available on GitHub [^3] https://github.com/The-Art-of-Story-Telling-2025/The-Art-of-Story-Telling-2025-The-Art-of-Story-Telling-2025.
@@ 2 System Architecture
Figure 1 shows the multi-agent architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy and Absentation) distributed throughout these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system’s front-end design is shown in Appendix B.
Once the five phases are completed, the input is passed to the Writer agent. The Writer generates a story based on the user’s inputs. The generated story is then reviewed by the Reviewer agent. The Reviewer checks whether the story is suitable for children, making any necessary edits to ensure it is age-appropriate.
After the story is finalized, it is sent to the Narrator agent, which converts the text into natural speech. Simultaneously, the story is also forwarded to the Film Director agent, who writes a detailed scene guide for each paragraph, outlining the visual elements needed to represent the narrative.
For each paragraph’s scene guide, the Music Director agent generates a music guide that reflects the emotional tone and context of the story. This guide is passed to the Musician agent, which creates the corresponding background music. Both the scene guide and the music are then used by the Animator agent to generate a video for a cohesive and immersive animation.
The final outputs include the textual story, the narrated audio, and the animated video with background music, creating a multimodal storytelling experience for children.
@@ 3.1 LLMs for Story generation
3.2 Story Evaluation
There is currently no established gold standard for evaluating stories, particularly those generated by large language models. However, different studies have used their own set of metrics to evaluate the quality of a story from a language, structure, and content perspective. Yang and Jin (2024) list a set of criteria that prior studies have used. These criteria are both story specific e.g. style and language, and LLM specific, e.g. naturalness and non-hallucination. This approach not only measures the quality of the generated text as a story but also checks for common errors associated with LLMs in generation tasks. Akoury et al. (2020) further relied on crowd-sourced evaluation and feedback from user interviews to gauge the quality of their stories. Other studies (Chhun et al., 2022) go beyond just the linguistic and structural aspects of the stories but also evaluate the content with metrics like narrative flow and emotions. Researchers from different groups have worked on evaluating the content of the stories from a “story-writing” perspective as well. Xie et al. (2023b) use metrics like fluency, coherence, and Interestingness to present a holistic evaluation framework. Another study by the same group Xie et al. (2023a) evaluates the stories by measuring the closeness between the generated stories and equivalent human-written stories
3.3 Evaluation of Text-to-Speech Systems
Both numerical and subjective metrics have been used to evaluate data produced by text-to-speech systems. Martín et al. (2020) use the mean opinion score across metrics like pronunciation, comprehension, etc., for multiple TTS systems. Another study Alvarez and Huckvale (2002) uses a similar metric-based method coupled with other measures like the Pair Comparison test where the listener is presented with two audios from different systems and chooses which one they prefer. Valentini-Botinhao et al. (2011) also use the pair comparison test in their study. Cooper et al. (2024) gives a detailed overview of objective and subjective metrics used for audio evaluation, and emphasizes the importance of metrics like "naturalness" and "intelligibility".
3.4 Evaluation of Text-to-Video Systems
Recently, there have been notable advancements in text-to-video models, and subsequently their evaluation. Zhang et al. (2024) detail a human evaluation protocol centered around metrics like video quality, temporal quality, and text alignment. Wu et al. (2024) introduce a numerical metric "Text-to-video" score calculated by metrics that measure both the text and video alignment and video quality. Hence, we can see both subjective and objective metrics being employed for evaluating synthesized videos.
4 Methodology
4.1 Evaluation
For any system that leverages generative AI for data generation, it is important to rely on a set of evaluation metrics to gauge the quality of the data. Our system uses three forms of data - story text generated through a LLM, audio narration of the stories generated through a TTS model, and the visuals for the stories generated using TTV models. Since each of these data is unique, we need a distinct set of evaluation metrics for each data type. The human evaluators for each module are six Computer Science students with English as the medium of education. Each criterion for LLM story generation, TTS, and TTV was evaluated by human raters, with scores assigned on a scale of 0, 1, or 2 (breakdown given in Appendix C).
4.1.1 LLMs for Story Generation
@@
Keeping prior literature and our interviews in mind, we are using the following framework to evaluate our stories.
@@
4.1.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, We gathered a diverse dataset of 100 stories from Projecgt Gutenberg, 50 appropriate and 50 inappropriate as rated by our evaluators.. Additionally, we generated a dataset of 50 stories using LLMs, which were similarly labeled as appropriate or inappropriate by human annotators. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix D.
4.1.3 Text-to-Speech Models
@@
4.1.4 Text-to-Video Models
For the TTV module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ‘animated‘ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the the scene guide. The system prompt for Director is given in Appendix D. The scene guide is then used for video generation. For each video, we assess the following metrics:
@@ 4.1.4 Text-to-Video Models
5 Results and Discussion
Throughout this section, the win rate reflects the proportion of times a model outperformed others, the tie rate indicates instances of comparable performance, and the loss rate captures the frequency of underperformance.
@@ 5.1 LLMs for Story Generation
Table 2 summarizes the pairwise comparisons among six models of varying scales—Gemma-2-9b, Gemma-2-27b (Google, 2024), Llama-3.1-8b, Llama-3.1-70b (Meta, 2024), GPT-4o, and GPT-4o-mini (OpenAI, 2024)—based on human evaluations of their story outputs.
A first notable observation is that simply scaling up parameter counts does not guarantee stronger performance in the downstream task of story generation. Gemma-2-9b outperforms its larger counterpart Gemma-2-27b (39.67% vs. 25.67% in win-rate), underscoring that bigger parameter counts are not always decisive for story-generation quality. Likewise Llama-3.1-70b achieves notable 43.33% wins and only 27.67% losses against GPT-4o. GPT-4o also performs poorly against GPT-4o-mini with only 16.33% win-rate. Across most pairwise comparisons, tie-rates hover near 30%, suggesting that many of these models are closely matched in generating coherent and engaging stories. Based on these head-to-head outcomes, Gemma-2-9b emerges as the most consistently strong performer overall, outperforming all the other models. Interestingly, the bigger model from the same family, Gemma-2-27b is the weakest based on the human evaluation, wining none of the pairwise comparisons.
@@
Overall, GPT-4o-mini emerges as most consistent across the board, consistently appearing at or near the top in most categories, especially Grammar, Appropriate Language, and Naturalness. Gemma-2-9b scores highest in Creativity and Structural Consistency, making it appealing for users. In contrast, GPT-4o sits at the lower end in several metrics (particularly Creativity and Structural Consistency), though it remains reasonably competitive in Grammar and Linguistic Consistency.
@@ 5.3 Text-to-Speech
Table 4 shows how the two text-to-speech models—XTTSv2 and StyleTTS 2—fared under female and male reference voices. When using the female reference audio, XTTSv2 achieves a 37.83% win-rate against StyleTTS 2 but also has a slightly higher loss-rate of 39.00%, reflecting a very close quality. By contrast, with the male reference audio, XTTSv2 performs somewhat better overall (39.67% win vs. 34.67% loss). Tie-rates remain in the 23–26% range for both scenarios, indicating that both models produce similarly acceptable outputs in a substantial fraction of cases.
Although these comparisons do not reveal a single dominant TTS system across all conditions, the trends suggest that XTTSv2 may slightly outperform StyleTTS 2 in handling the male voice, whereas StyleTTS 2 holds a marginal edge—or at least near parity—when cloning the female voice.
@@ 5.4 Text-to-Video
Table 5 summarizes the relative performance of three text-to-video animation styles—Cartoon, Anime, and Animated—based on head-to-head comparisons. Cartoon lags behind Anime (28.67% win vs. 48.33% loss) and also against Animated (28.33% win vs. 49.50% loss), suggesting that its visuals are generally of lower quality. On the other hand, Animated leads when paired with Anime (34.33% vs. 36.00% loss), although its margin is not overwhelmingly large. Tie-rates hover around 20–30%, reflecting that all three styles can produce some visually appealing results under certain prompts. Overall, Animated style appears slightly stronger than both Cartoon and Anime in crafting engaging visuals, whereas Cartoon tends to be the weakest performer in pairwise comparisons.

### v6

# Kahaani: A Multimodal Co-Creative Storytelling System

@@ Abstract
This paper introduces Kahaani, a multimodal, co-creative storytelling system that leverages Generative Artificial Intelligence, designed for children to address the challenge of sustaining engagement to foster educational narrative experiences. Here we define co-creative as a collaborative creative process in which both the child and Kahaani contribute to the generation of the story. The system combines Large Language Model (LLM), Text-to-Speech (TTS), Text-to-Music (TTM), and Text-to-Video (TTV) generation to produce a rich, immersive, and accessible storytelling experience. The system grounds the co-creation process in two classical storytelling framework, Freytag’s Pyramid and Propp’s Narrative Functions. The main goals of Kahaani are: (1) to help children improve their English skills, (2) to teach important life lessons through story morals, and (3) to help them understand how stories are structured, all in a fun and engaging way. We present evaluations for each AI component used, along with a user study involving three parent–child pairs to assess the overall experience and educational value of the system.
Kahaani: A Multimodal Co-Creative Storytelling System
Samee Arif1, *Muhammad Saad Haroon1, *Aamina Jamal Khan1, Taimoor Arif2, Agha Ali Raza1, Awais Athar3 1Lahore University of Management Sciences, 2University of Michigan, 3Strategize Labs {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@strategize.inc
@@ 1 Introduction
In this paper, we introduce Kahaani [^1] https://github.com/SaadH-077/kahaani, a system designed to co-create multimodal stories for children. The central problem Kahaani aims to address is the difficulty children often have in maintaining attention and developing narrative competence with traditional story delivery. To support children in overcoming these challenges, the system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^2] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^3] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases, exposition, rising action, climax, falling action, and resolution, ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills (Ciğerci and Yıldırım, 2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.
@@
The conversion of written text into spoken words provides auditory stimuli that can improve reading comprehension, especially for students with learning disabilities. Studies have shown that TTS can support struggling readers by offering audio input as digital text is read aloud, aiding in comprehension and retention (Keelor et al., 2020). Visual stimuli can enhance understanding and retention of information. Research indicates that combining text and images in multimedia explanations is more effective than text alone, as it caters to diverse learning preferences and helps maintain attention (Liu, 2024). Background music adds an emotional layer to narratives, facilitating deeper engagement and supporting memory formation. Music has been recognized for its ability to evoke emotions and create a captivating environment, which can enhance memory retention by increasing engagement (Sinoradzki, 2023). By integrating these components, our system addresses the limitations of traditional storytelling methods that often lack multimodal stimuli, leading to reduced attention spans and engagement. This multimodal approach ensures that storytelling is not only entertaining but also educational, catering to various learning styles and promoting cognitive development.
@@ 2 System Architecture
Figure 2 shows the architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy and Absentation) distributed throughout these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system’s input design is shown in Figure 1.
@@
Once the five phases are completed, the input is passed to the Writer. The Writer LLM generates a story based on the user’s inputs. The generated story is then reviewed by the Reviewer LLM. The Reviewer checks whether the story is suitable for children, making any necessary edits to ensure it is age-appropriate. After the story is finalized, it is sent to the Narrator (TTS), which converts the text into natural speech. Simultaneously, the story is also forwarded to the Film Director LLM, which writes a detailed scene guide for each paragraph, outlining the visual elements needed to represent the narrative. For each paragraph’s scene guide, the Music Director LLM generates a music guide that reflects the emotional tone and context of the story. This guide is passed to the Musician (TTM), which creates the corresponding background music. Both the scene guide and the music are then used by the Animator (TTV) to generate a video for a cohesive and immersive animation. The final outputs include the textual story, the narrated audio, and the animated video with background music, creating a multimodal storytelling experience for children.
@@ 3 Related Work
While existing story-generation systems focus primarily on text-based plot development and iterative planning, they largely omit multimodal, co-creative frameworks that integrate narrative text with audio, animation, and music for educational and engagement-focused applications, which is the gap Kahaani aims to fill.
4 Experimental Setup
We conduct human evaluation of LLM for story generation, TTS narration, TTV animation. The human evaluators for each module were six Computer Science students with English as the medium of education. Each criterion was assigned a score of of 0, 1, or 2 (breakdown given in Appendix Section C). After evaluating the components individually, we selected the top-performing models for each of the other modules and used them to conduct the user study of the complete system.
@@ 4.1 LLMs for Story Generation
In order to understand how teachers evaluate stories submitted by their students, we conducted semi-structured interviews with 3 teachers who teach English as a subject to students of age group 6 to 12. The teachers emphasized assessing the content of the story for its quality. This includes evaluating the language used, checking the logical sequence of the story, usage of “key terms” to “cue the reader to the direction of the discourse”. The teachers noted:
“…..we try not to use any harmful/abusive language”
“We check if there is a clear beginning and end”
“Does the writer stick to the topic of the story?”
Other than the content, the structure of the story is also scrutinized. One of the teachers said:
“A group of words does not make a complete sentence. Sentences that are made up of a group of words without a direction need to be penalized”
Keeping prior literature and our interviews in mind, we use the following framework to evaluate our stories.
@@
4.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, We gathered a diverse dataset of 100 stories from Project Gutenberg, 50 appropriate and 50 inappropriate as rated by our evaluators. Additionally, we generated a dataset of 50 stories using LLMs, which were similarly labeled as appropriate or inappropriate. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix Section D.
4.3 Text-to-Speech Models
@@
4.4 Text-to-Video Models
For the TTV module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ‘animated‘ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the scene guide. The system prompt for Director is given in Appendix D. The scene guide is then used for video generation. For each video, we assess the following metrics:
@@ 4.4 Text-to-Video Models
4.5 User Study
To evaluate the overall experience and educational value of the system, we conducted a user study with three parent-child pairs. Each child interacted with the system to co-create a story, and both children and their parents were asked to provide feedback through structured questionnaires. The goal was to assess the system’s usability, engagement, age-appropriateness, and educational impact from both perspectives. To make it easier for children to respond, we included mostly rating-based questions (represented with a * below) on a 1–5 scale.
Children’s Questionnaire
  1. 1.
    How easy was the story to understand?*
  2. 2.
    How likely are you to recommend this system to your friends?*
  3. 3.
    How much did you like the animations?*
  4. 4.
    How much did you like the way the story was told (e.g., narration, voice)?*
  5. 5.
    How much would you rate the generated story itself?*
  6. 6.
    Overall, how would you rate your experience with the storytelling system?*
  7. 7.
    Did you learn something new from the story? What was it?
  8. 8.
    What changes would you suggest to improve the overall experience?
Parents’ Questionnaire
  1. 1.
    How appropriate was the content for your child’s age?*
  2. 2.
    How engaged was your child while using the system?*
  3. 3.
    How likely are you to recommend this system to other parents?*
  4. 4.
    How much did you like the system’s design (e.g., visuals, ease of use)?*
  5. 5.
    Overall, how satisfied are you with the storytelling system?*
  6. 6.
    How did your child react to using the storytelling system?
  7. 7.
    Do you think the system helped your child learn something new?
  8. 8.
    Do you think this system will have any improvement in your child’s language skills or creativity?
5 Results & Discussion
@@ 5.1 LLMs for Story Generation
Table 2 summarizes the pairwise comparisons among Gemma-2-9b, Gemma-2-27b (Google, 2024), Llama-3.1-8b, Llama-3.1-70b (Meta, 2024), GPT-4o, and GPT-4o-mini (OpenAI, 2024) based on human evaluations of their story outputs.
@@
A notable observation is that simply scaling up parameter counts does not guarantee stronger performance in the downstream task of story generation. Gemma-2-9b outperforms its larger counterpart Gemma-2-27b (39.67% vs. 25.67% in win-rate), underscoring that bigger parameter counts are not always decisive for story-generation quality. Likewise Llama-3.1-70b achieves notable 43.33% wins and only 27.67% losses against GPT-4o. GPT-4o also performs poorly against GPT-4o-mini with only 16.33% win-rate. Across most pairwise comparisons, tie-rates hover near 30%, suggesting that many of these models are closely matched in generating coherent and engaging stories. Based on these head-to-head outcomes, Gemma-2-9b emerges as the most consistently strong performer overall, outperforming all the other models. Interestingly, the bigger model from the same family, Gemma-2-27b is the weakest based on the human evaluation, wining none of the pairwise comparisons.
In Table 3, we use the Bradley–Terry model (Firth, 2005) to rank the LLMs based on their pairwise comparison results, with ties counted as half-wins. The top-performing model for our generative task is Gemma-9b, followed by GPT-4o-mini. Interestingly, for both Gemma and GPT, the smaller models outperform their larger counterparts, whereas for Llama, the larger Llama-70b slightly outperforms the smaller Llama-8b. Overall, these results suggest that smaller models tend to excel at creative and linguistic tasks. While larger models offer longer context windows and improved accuracy, this has a tradeoff against long-form, creative generation tasks.
Rank Model BT Strength ($\pi_{i}$)
1 Gemma-9b 1.224
2 GPT-4o-mini 1.097
3 Llama-70b 1.058
4 Llama-8b 0.983
5 Gemma-27b 0.884
6 GPT-4o 0.810
Table 3: Bradley-Terry (Firth, 2005) ranking of models based on pairwise comparisons, with ties counted as half-wins. Higher $\pi_{i}$ indicates a stronger model.
@@
Overall, GPT-4o-mini emerges as most consistent across the board, consistently appearing at or near the top in most categories, especially Grammar, Appropriate Language, and Naturalness. Gemma-2-9b scores highest in Creativity and Structural Consistency, making it appealing for users. In contrast, GPT-4o sits at the lower end in several metrics particularly Creativity and Structural Consistency.
@@ 5.3 Text-to-Speech
Table 5 shows how the two text-to-speech models, XTTSv2 and StyleTTS 2, fared under female and male reference voices. When using the female reference audio, XTTSv2 achieves a 37.83% win-rate against StyleTTS 2 but also has a slightly higher loss-rate of 39.00%, reflecting a very close quality. By contrast, with the male reference audio, XTTSv2 performs somewhat better overall (39.67% win vs. 34.67% loss). Tie-rates remain in the 23–26% range for both scenarios, indicating that both models produce similarly acceptable outputs in a substantial fraction of cases. Although these comparisons do not reveal a single dominant TTS system across all conditions, the trends suggest that XTTSv2 may slightly outperform StyleTTS 2 in handling the male voice, whereas StyleTTS 2 holds a marginal edge when cloning the female voice.
@@ 5.4 Text-to-Video
Table 6 summarizes the relative performance of three text-to-video animation styles—Cartoon, Anime, and Animated—based on head-to-head comparisons. Cartoon lags behind Anime (28.67% win vs. 48.33% loss) and also against Animated (28.33% win vs. 49.50% loss), suggesting that its visuals are generally of lower quality. On the other hand, Animated leads when paired with Anime (34.33% vs. 36.00% loss), although its margin is not overwhelmingly large. Tie-rates hover around 20–30%, reflecting that all three styles can produce some visually appealing results under certain prompts.
@@
We also ranked these three models in Table 7 using the Bradley-Terry model (Firth, 2005), with Animated winning as the top ranked model, closely followed by Anime. Overall, Animated style appears slightly stronger in crafting engaging visuals, whereas Cartoon tends to be the weakest performer.
Rank Style BT Strength ($\pi_{i}$)
1 Animated 1.167
2 Anime 1.129
3 Cartoon 0.759
Table 7: Bradley-Terry (Firth, 2005) ranking of animation styles based on pairwise comparisons, with ties counted as half-wins. Higher $\pi_{i}$ indicates a stronger or more preferred style.